Given: IA = 2 mA, ID = - 0.1 mA, IG = IH = IK

(a) Find the numerical value of IB in mA.

(b) Find the numerical value of IH in mA.

(c) If VA = - 3 V and VG = 2 V, find the numerical value of the power in element B in mW and whether it is supplying or receiving power.

(d) If also VF = - 100 mV, what is the power in element H in mW?

(e) If also VE = - 150 mV, what is the power in element C and D together in mW?
Answer #1

(a) Element A and B lie in same branch of the circuit. Therefore, the magnitude of the current is same in both elements. The-V+V-VG=0 V, =Va+V = 2 V+3 V =5 V Therefore, the magnitude of V, is 5 V. The current through the element is -2 mA. Therefore,|PH=\V#1: =(2,1V)(0.63 mA)(1 = 13 W